6. Two cars are traveling along a straight road. Car (A) maintains a constant speed of 780 km/h; car (B) maintains a constant speed of 104 km/h. Att o, car (B) is 430 km behind car A. How far will car (A) travel from t O before it is overtaken by car (B)? km
Answer #1


for car A


distance travelled in time t dA = vA*t = 78*t

for car B


distance travelled in time t dB = vB*t = 104*t


dB - dA = 43 km

104*t - 78*t = 43

t = 43/26 h


distance travelled by car A = dA = vA*t = 78*43/26 = 129 km